Daytime AO work on 4 June 2011

Observer(s): D. Miller, L Busoni, A. Puglisi, S. Esposito (LBTO)

Telescope Operator: DHuerta (LBTO)

SW Support: (Tucson)

Telescope Support: J. Urban (LBTO)

Instrument Support: (LBTO)

Summary

The retroreflector was not removed last night because OSU was observing with LUCI. Thus, the telescope and retroreflector were ready to go when we woke up this morning.

Simone and Doug started up the AO system at about 8:30. Doug sent Alfio and Luca emails with a couple of changes to the default startup state of the WFS.

URTH fans were turned off. Tests of image quality with fans on and off are scheduled for today, as well as testing and debugging of non-common path aberration routines. During the Italian's lunch (12:30 ish) Doug will test his final (hopefully) versions of the IDL routines to communicate with PISCES and fits file create (header info and corquad) routines.

Enrico is scheduled to arrive on the mountain with Marco Pedani this afternoon.

10:00 Doug created a PISCES Setup Instructions page

PISCES Setup Instructions

10:30 The routines to automatically determine non-common path aberrations are working well and an improved strehl for the FeII filter (80ish to 90ish) in diffraction limited mode (no atmospheric turbulence) was determine in about 30 minutes. Now to the H filter. This will be a good test of these N.C.P. routines.

11:00 Doug submitted IT #3425 and #3426 to Sys. Admin. to added Pisces to the DNS server and a warning when we ssh into wfsdx

While the Italians were ate lunch, Doug tested his PISCES routines. All the "low level" routines needed to set parameters and take images on PISCES have been tested and debugged. The "high level" routine have been tested (eg pisces_get_image.pro to start an image "go" and then poll until the file is written in /newdata, run corquad on the saved, read TCS DD variable, add these variables to the header and write out a new file.

The H filter images still had problems and the application of automatic slope null routine did on help. It's still a mystery.

When hand over started, Doug shut off the dry air to the PISCES window. He and Joe K also connected the fiber from the DX M2 Jbox to the patch panel in the ULTH, to fiber that runs to the 2nd floor computer room switch. Joe had tested the fiber from the ULTH to the 2nd floor, but has not tested from the ULTH to the DX M2 Jbox. There is no reason to suspect this fiber, but in case there is a problem with connection, it could be this upper fiber. The fiber, IP and switch should now be ready for the second Anemometer.

Jeff moved the telescope to horizon at about 4:30 to remove the retroreflector.

Doug left for Tucson at 4:45 and handed over the truck to Guido so he can drive to the LBTO tomorrow.

-- DougMiller - 04 Jun 2011
Topic revision: r3 - 06 Jun 2011, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback